Suppose the Price/Earnings Ratio for the S&P 500 is 15 and the dividend payout ratio of...

  1. Suppose the Price/Earnings Ratio for the S&P 500 is 15 and the dividend payout ratio of the S&P 500 is 28%. The future growth rate of dividends is expected to be 3.35%.
    1. Compute the expected return of the Market.
    2. Use Goal Seek or Solver to determine the dividend growth rate that would yield an expected Market return of 7%.
Answer #1

a.

Using Constant Growth Model,

r = D1/Price + g

r = E1(0.28)/Price + 0.0335

r = 0.28/15 + 0.0335

r = 5.22%

b.

0.07 = 0.28/15 + g

g = 5.13%